# Krijgslaan, 281-S9, 9000 Gent, Belgium

### v1

@@ 3.1 An exactly solvable $U(1)^{3}$ model for the radiation
In this section we briefly introduce a specific model for the radiation part of the action, for which the complete set of equations (including the ones for the gauge fields) can be solved exactly. This is a model with a $U(1)^{3}$ symmetry (or in other words containing a triplet of gauge fields $A_{\mu}^{(I)},\,I=1,2,3$). It was first studied in a similar context in Marolf:2021kjc (for a fixed negative cosmological constant), but here we consider its extension, since we include the presence of the non-trivial inflaton field $\phi$, see also Lan:2024gnv . We shall also improve the previous analyses of this model using differential forms for the gauge field, allowing us to define the solutions in a more invariant manner.
@@
and the isotropic ansatz that we consider [^10] This is invariant under the left action of $SU(2)$ and hence isotropic on $S^{3}$. A similar ansatz for an $SU(2)$ non-Abelian gauge field gives rise to the Meron wormhole, see Maldacena:2004rf ; Betzios:2019rds and references within. See also Rey:1989th for a similar ansatz and manipulations, in the non-abelian case. is
@@
where $A^{-}_{\mu},~A^{+}_{\mu}$ correspond to the two possible modes at the boundary. For a massless gauge field the conformal dimensions are given by $\Delta_{\pm}=\frac{d}{2}\pm\left(\frac{d}{2}-1\right)\Rightarrow\Delta_{+}=2\,,\,\Delta_{-}=1$ where we specialize to $d=3$. This is in agreement with the values of $\lambda_{\pm}$ found just above since $\lambda_{\pm}=\Delta_{\pm}-1$. This simple check clarifies that the leading asymptotic behavior of the electromagnetic potential is a radially independent constant corresponding to a boundary source using the standard Dirichlet BCs. With these conditions the $U(1)^{3}$ symmetry at the boundary is a global symmetry and the magnetic field approaches a source at the boundary (to leading order) since $B_{I}\sim H^{(0)}$, while the electric field only contains a VEV part (since $E_{I}\sim H^{\prime}(\tau)\sim H^{(1)}$). Notice that very importantly, in three boundary dimensions both leading and subleading modes are normalizable, since $\Delta_{-}=1<d/2=3/2$ (and $d>2)$. This means that we can also interpret the leading behaviour as a VEV and the subleading as a source (alternate quantization/Neumann BCs). In this later case the $U(1)^{3}$ is a local gauge symmetry on the boundary. The transformation between these two options is affected by a boundary Legendre transform. This transformation is also one part/generator of the $SL(2,Z)$ group action that performs Electric/Magnetic (E/M) duality in the bulk. In other words E/M duality contains both the Legendre transform together with a Hodge duality operation that exchanges the E/M fields Hawking:1995ap ; Witten:2003ya  [^11] In fact one is also free to add a Cherns-Simons action $S_{C.S.}\sim k\int A\wedge dA$ at the boundary and the T-operation of E/M duality ($SL(2,Z)$) shifts the C.S. level $k$ by one unit.. In order to perform the change from Dirichlet to Neumann BCs (alternate quantization), one needs to add the following boundary action (for a single boundary component [^12] In the case of wormholes (two boundary components), provided the wormhole is $\mathbb{Z}_{2}$-symmetric at $\tau=0$ ($a$ even) and $H(\tau)$ is even, the contribution is simply doubled.)
@@ 3.2 Background subtraction/renormalization procedure in EFLRW metric
  • Holographic renormalization Skenderis:2002wp : add local covariant counterterms on a cut-off surface to render each on-shell action finite. The resulting renormalized action is scheme dependent, with the minimal subtraction scheme adding only the minimal terms required to cancel divergences.
@@
Although the renormalized on-shell action in holographic renormalization depends on the choice of finite counterterms i.e. choice of “scheme”, differences between saddles evaluated with the same boundary conditions are well defined and scheme independent. In Marolf:2021kjc the authors used holographic renormalization; here we instead implement background subtraction (but we still analyze the appropriate counterterms and renormalized action for the scalar field since they help clarify the physics and the phase space of solutions).
@@
with $V_{\rm AdS}<0$ and $m^{2}<0$. Thus $\phi=0$ is an extremum of the potential and always solves the scalar equation of motion (cf. the first line of (9)). The $\phi=0$ saddles may equivalently be viewed as solutions of Einstein gravity with a negative cosmological constant $\Lambda=V_{\rm AdS}$; these are precisely the configurations considered in Marolf:2021kjc .
@@ 4 A menagerie of (Euclidean) gravitational saddles
  • More general oscillatory wormholes (see Jonas:2023ipa for some analogous asymptotically flat geometries), for which the scale factor oscillates in the throat region multiple times. Such geometries can be found in our class of models iff the scalar potential is also an oscillating function in the region where it is positive (we term this a quasi-oscillatory behaviour).
@@
The first type of geometries does not contribute to cosmological transition amplitudes (only affects the overall normalization of probabilities, see our discussion in section 2) and concerns the probability that the system will remain in an AdS (non-cosmological) background. The second type of geometries leads to crunching cosmological solutions upon rotating to Lorentzian signature, see Maldacena:2004rf ; Betzios:2019rds ; VanRaamsdonk:2021qgv ; Betzios:2024oli for discussions. The third type of backgrounds leads to expanding/inflating cosmologies Betzios:2024oli ; Betzios:2024zhf and is the one relevant for phenomenological purposes. Expanding cosmologies can also be produced by the oscillatory wormholes, if the number of oscillations is odd, but these require (quasi)-periodic potentials with further non-trivial features as we shall prove.
Finally in the literature there are also appearances of single boundary “centaur” geometries Anninos:2017hhn , which can be thought of as a finely tuned case of the wineglass wormholes, when the second asymptotic region decouples and the manifold smoothly caps off. These configurations are not $\mathbb{Z}_{2}$ symmetric in the radial direction and in our setting they are expected to exist when the running scalar reaches a local minimum/maximum of the scalar potential with a positive value and stops there. We find that in our class of models, the presence of the flux that supports the minimum size of the throat, is in clash with a subsequent smooth closure of the Euclidean manifold. In other words such solutions can only be found if the stress-energy density becomes a time dependent function that while being non-trivial initially, it diminishes as the scalar field reaches the local maximum/minimum of the potential, leading to a pure dS solution in the deep interior of the geometry.
@@ 4.3 Wineglass wormhole solutions with running scalar
In a more general setting, the scalar/inflaton potential being a non-trivial function, leads to (Holographic) RG flows. The scalar becomes then a non-trivial function of the radial coordinate $\phi(\tau)$ , reaching negatively valued local minima/maxima of the potential in the asymptotic AdS boundaries. In the present setting, we shall be interested in flows for which the background geometry resembles a “wineglass wormhole” [^13] This can also be thought of as a Euclidean version of a “bag of gold” space (see Fu:2019oyc and references within). While in the Lorentzian case the “bag of gold” is a codimension one surface, here the complete Euclidean spacetime has the form of a “bag of gold”, see fig. 1.. On such backgrounds the scalar explores a positive region of the potential, near the middle of the wormhole throat, before returning back to an AdS negative local maximum. In general it is difficult to find a completely analytic solution of the Euclidean saddle point EOM. Here we shall resort to a smooth piecewise (continuous and thrice differentiable) ansatz, to overcome this obstacle and yield the solutions of interest.
@@ 4.3.1 Piecewise wineglass wormhole solutions
In figs. 3 we depict the scale factor (top panel) and its first (bottom left) and second derivative (bottom right) and in figs. 4 the potential as a function of $\phi$ (top panel) and $\tau$ (bottom left) and the scalar field value as a function of $\tau$ (bottom right), for $\tau\leq 0$. The behaviour for $\tau>0$ is identical by symmetry, since all functions involved are even in $\tau$. We observe that both the potential and the running of the scalar are consistent with the criteria laid out in Betzios:2024oli ; Betzios:2024zhf for the existence of Euclidean wineglass wormhole solutions and that all the analytic functions are smoothly joined in the interpolating region, up to third order in derivatives. The parameter choice used in figs. 3 and 4 is $H_{\rm dS}=1$, $H_{\rm AdS}=1.1$, $A_{\rm dS}=1$ and $c_{\rm AdS}=-2\sqrt{A_{\rm AdS}(A_{\rm AdS}-A_{\rm dS})}=-2$ , which corresponds to the values of the physical parameters $\kappa\tilde{V}_{0}\simeq 1.37$, $\kappa\tilde{V}_{\rm AdS}=-1.21$, and the derived parameter $\kappa\tilde{V}_{min}\simeq-8.97$.
@@
where the lower bound is the BF-bound Breitenlohner:1982bm ; Breitenlohner:1982jf . Moreover, since $\Delta_{-}=1<d/2=3/2$ , we are in a situation where we can choose multiple different quantizations for the scalar field, since both modes are normalizable. In particular if we expand it around infinity we find that
@@
Two typical choices correspond to setting the coefficient of the $\Delta_{-}$ term to be the source and that of $\Delta_{+}$ to be the VEV (Dirichlet BCs), and the other is the case in which $\tilde{\phi}_{-}$ corresponds to the VEV and $\tilde{\phi}_{+}$ to the source (Neumann BCs) [^16] In general one can consider mixed BCs leading to generic multi-trace deformations of the boundary CFT.. In addition, we observe that we are currently in a resonant setup, in the sense that we also have $\Delta_{+}-\Delta_{-}=1$ , so that the first subleading term in the expansion of $\tilde{\phi}_{-}$ scales the same as the leading term in the expansion of $\tilde{\phi}_{+}$ and so forth [^17] Nevertheless, no logarithmic term (associated with a conformal anomaly) appears in the expansion, since the conformal dimension satisfies $\Delta_{+}\notin\frac{d}{2}+\mathbb{N}$ (see the criterion in Petkou_1999 )..
@@
We then write down the scalar counterterm action (for a single boundary component) [^18] This counterterm action is the same for all different choices of quantization. In general it is derived integrating the relation $\delta S_{c.t.}(\tilde{\phi})=\int_{\partial\mathcal{M}}d^{3}x\sqrt{h}\sum_{n<\Delta_{+}}\tilde{\pi}_{n}\delta\tilde{\phi}$ , where $\tilde{\pi}_{n}$ are the various terms in the expansion of the momentum., see Papadimitriou:2007sj and references within
@@
In our specific case where $\Delta_{+}=2,\,\Delta_{-}=1<3/2$, only the first term in this action is needed for renormalization Papadimitriou:2007sj . We then compute the counterterm contribution to the momentum from (70) to be $\tilde{\Pi}_{c.t.}=\sqrt{h}\Delta_{-}H_{\rm AdS}\tilde{\phi}$. Subsequently, the renormalized momentum is given by the sum of the bare momentum (69) and the counterterm $\tilde{\Pi}_{c.t.}$ i.e.
@@ 4.3.2 The wineglass E/M gauge field background
Furthermore, if we wish to understand the wormhole background as providing good initial conditions for phenomenologically realistic inflation as in the scenario proposed in Betzios:2024oli ; Betzios:2024zhf , we can constrain the parameters of the model further. In particular assuming that at $\tau=t=0$, i.e. at the onset of standard Lorentzian evolution, the field is already located in the plateau region we obtain that,
$\tilde{V}_{0}\lesssim 3.73\times 10^{-10}\kappa^{-1}\left(\frac{r}{0.036}\right)\,$, (86)
where $r$ is the tensor-to-scalar ratio, bounded by $r<0.036$ at $95\%$ CL when combining the latest Planck, BICEP/Keck, and BAO data Planck:2018jri ; BICEP:2021xfz . Equations (56) and (57) allow one to relate the value of the scalar potential at $\tau=0$ to that of the E/M sector, namely
@@
$|H_{0,WG}|\gtrsim 3.17\times 10^{4}\left(\frac{0.036}{r}\right)^{1/2}\,.$ (88)
@@ 4.4.1 Piecewise oscillatory wormhole solutions
These results for the (quasi)-oscillatory wormholes are very important and contain crucial differences and resolve some paradoxes that appear in analogous (exactly) periodic solutions, such as the axion de-Sitter wormholes of Aguilar-Gutierrez:2023ril . The axion de-Sitter wormholes exist for a fixed cosmological constant, have no UV AdS boundaries and give rise to a serious paradox, since their number of periods is not physically bounded by any principle. Moreover, in that case each period lowers the on-shell action, rendering it unbounded from below in the limit of infinite periods [^21] In a sense this issue is related to the well-known conformal mode problem of Euclidean gravity, exacerbated by the shift symmetry of the constant potential (due to the fixed cosmological constant), see also Blommaert:2025bgd .. Here the presence of the AdS boundaries and the fact that the potential is a non trivial (quasi)-periodic function whose number of periods is directly correlated to the number of geometric periods in the IR throat part of the geometry, precludes such issues.

### v2

@@ 3.1 An exactly solvable $U(1)^{3}$ model for the radiation
In this section we briefly introduce a specific model for the radiation part of the action, for which the complete set of equations (including the ones for the gauge fields) can be solved exactly. This is a model with a $U(1)^{3}$ symmetry (or in other words containing a triplet of gauge fields $A_{\mu}^{(I)},\,I=1,2,3$). It was first studied in a similar context in Marolf and Santos (2021) (for a fixed negative cosmological constant), but here we consider its extension, since we include the presence of the non-trivial inflaton field $\phi$, see also Lan and Piao (2024). We shall also improve the previous analyses of this model using differential forms for the gauge field, allowing us to define the solutions in a more invariant manner.
@@
and the isotropic ansatz that we consider [^10] This is invariant under the left action of $SU(2)$ and hence isotropic on $S^{3}$. A similar ansatz for an $SU(2)$ non-Abelian gauge field gives rise to the Meron wormhole, see Maldacena and Maoz (2004); Betzios et al. (2019) and references within. See also Rey (1990) for a similar ansatz and manipulations, in the non-abelian case. is
@@
where $A^{-}_{\mu},~A^{+}_{\mu}$ correspond to the two possible modes at the boundary. For a massless gauge field the conformal dimensions are given by $\Delta_{\pm}=\frac{d}{2}\pm\left(\frac{d}{2}-1\right)\Rightarrow\Delta_{+}=2\,,\,\Delta_{-}=1$ where we specialize to $d=3$. This is in agreement with the values of $\lambda_{\pm}$ found just above since $\lambda_{\pm}=\Delta_{\pm}-1$. This simple check clarifies that the leading asymptotic behavior of the electromagnetic potential is a radially independent constant corresponding to a boundary source using the standard Dirichlet BCs. With these conditions the $U(1)^{3}$ symmetry at the boundary is a global symmetry and the magnetic field approaches a source at the boundary (to leading order) since $B_{I}\sim H^{(0)}$, while the electric field only contains a VEV part (since $E_{I}\sim H^{\prime}(\tau)\sim H^{(1)}$). Notice that very importantly, in three boundary dimensions both leading and subleading modes are normalizable, since $\Delta_{-}=1<d/2=3/2$ (and $d>2)$. This means that we can also interpret the leading behaviour as a VEV and the subleading as a source (alternate quantization/Neumann BCs). In this later case the $U(1)^{3}$ is a local gauge symmetry on the boundary. The transformation between these two options is affected by a boundary Legendre transform. This transformation is also one part/generator of the $SL(2,Z)$ group action that performs Electric/Magnetic (E/M) duality in the bulk. In other words E/M duality contains both the Legendre transform together with a Hodge duality operation that exchanges the E/M fields Hawking and Ross (1995); Witten (2003) [^11] In fact one is also free to add a Cherns-Simons action $S_{C.S.}\sim k\int A\wedge dA$ at the boundary and the T-operation of E/M duality ($SL(2,Z)$) shifts the C.S. level $k$ by one unit.. In order to perform the change from Dirichlet to Neumann BCs (alternate quantization), one needs to add the following boundary action (for a single boundary component [^12] In the case of wormholes (two boundary components), provided the wormhole is $\mathbb{Z}_{2}$-symmetric at $\tau=0$ ($a$ even) and $H(\tau)$ is even, the contribution is simply doubled.)
@@ 3.2 Background subtraction/renormalization procedure in EFLRW metric
  • Holographic renormalization Skenderis (2002): add local covariant counterterms on a cut-off surface to render each on-shell action finite. The resulting renormalized action is scheme dependent, with the minimal subtraction scheme adding only the minimal terms required to cancel divergences.
@@
Although the renormalized on-shell action in holographic renormalization depends on the choice of finite counterterms i.e. choice of “scheme”, differences between saddles evaluated with the same boundary conditions are well defined and scheme independent. In Marolf and Santos (2021) the authors used holographic renormalization; here we instead implement background subtraction (but we still analyze the appropriate counterterms and renormalized action for the scalar field since they help clarify the physics and the phase space of solutions).
@@
with $V_{\rm AdS}<0$ and $m^{2}<0$. Thus $\phi=0$ is an extremum of the potential and always solves the scalar equation of motion (cf. the first line of (9)). The $\phi=0$ saddles may equivalently be viewed as solutions of Einstein gravity with a negative cosmological constant $\Lambda=V_{\rm AdS}$; these are precisely the configurations considered in Marolf and Santos (2021).
@@ 4 A menagerie of (Euclidean) gravitational saddles
  • More general oscillatory wormholes (see Jonas et al. (2023) for some analogous asymptotically flat geometries), for which the scale factor oscillates in the throat region multiple times. Such geometries can be found in our class of models iff the scalar potential is also an oscillating function in the region where it is positive (we term this a quasi-oscillatory behaviour) [^13] There is also a fine-tuned case, whereby the potential is a fixed positive cosmological constant. This case admits exactly oscillatory solutions with a potentially troublesome infinite number of oscillations, such as in the axion de-Sitter wormholes studied in Halliwell and Myers (1989); Aguilar-Gutierrez et al. (2023). Lifting the flat direction of the potential precludes this possibility..
@@
The first type of geometries does not contribute to cosmological transition amplitudes (only affects the overall normalization of probabilities, see our discussion in section 2) and concerns the probability that the system will remain in an AdS (non-cosmological) background. The second type of geometries leads to crunching cosmological solutions upon rotating to Lorentzian signature, see Maldacena and Maoz (2004); Betzios et al. (2019); Van Raamsdonk (2022); Betzios and Papadoulaki (2024a) for discussions. The third type of backgrounds leads to expanding/inflating cosmologies Betzios and Papadoulaki (2024a); Betzios et al. (2025) and is the one relevant for phenomenological purposes. Expanding cosmologies can also be produced by the oscillatory wormholes, if the number of oscillations is odd, but these require (quasi)-periodic potentials with further non-trivial features as we shall prove.
Finally in the literature there are also appearances of single boundary “centaur” geometries Anninos and Hofman (2018), which can be thought of as a finely tuned case of the wineglass wormholes, when the second asymptotic region decouples and the manifold smoothly caps off. These configurations are not $\mathbb{Z}_{2}$ symmetric in the radial direction and in our setting they are expected to exist when the running scalar reaches a local minimum/maximum of the scalar potential with a positive value and stops there. We find that in our class of models, the presence of the flux that supports the minimum size of the throat, is in clash with a subsequent smooth closure of the Euclidean manifold. In other words such solutions can only be found if the stress-energy density becomes a time dependent function that while being non-trivial initially, it diminishes as the scalar field reaches the local maximum/minimum of the potential, leading to a pure dS solution in the deep interior of the geometry.
@@ 4.3 Wineglass wormhole solutions with running scalar
In a more general setting, the scalar/inflaton potential being a non-trivial function, leads to (Holographic) RG flows. The scalar becomes then a non-trivial function of the radial coordinate $\phi(\tau)$ , reaching negatively valued local minima/maxima of the potential in the asymptotic AdS boundaries. In the present setting, we shall be interested in flows for which the background geometry resembles a “wineglass wormhole” [^14] This can also be thought of as a Euclidean version of a “bag of gold” space (see Fu and Marolf (2019) and references within). While in the Lorentzian case the “bag of gold” is a codimension one surface, here the complete Euclidean spacetime has the form of a “bag of gold”, see fig. 1.. On such backgrounds the scalar explores a positive region of the potential, near the middle of the wormhole throat, before returning back to an AdS negative local maximum. In general it is difficult to find a completely analytic solution of the Euclidean saddle point EOM. Here we shall resort to a smooth piecewise (continuous and thrice differentiable) ansatz, to overcome this obstacle and yield the solutions of interest.
@@ 4.3.1 Piecewise wineglass wormhole solutions
In figs. 3 we depict the scale factor (top panel) and its first (bottom left) and second derivative (bottom right) and in figs. 4 the potential as a function of $\phi$ (top panel) and $\tau$ (bottom left) and the scalar field value as a function of $\tau$ (bottom right), for $\tau\leq 0$. The behaviour for $\tau>0$ is identical by symmetry, since all functions involved are even in $\tau$. We observe that both the potential and the running of the scalar are consistent with the criteria laid out in Betzios and Papadoulaki (2024a); Betzios et al. (2025) for the existence of Euclidean wineglass wormhole solutions and that all the analytic functions are smoothly joined in the interpolating region, up to third order in derivatives. The parameter choice used in figs. 3 and 4 is $H_{\rm dS}=1$, $H_{\rm AdS}=1.1$, $A_{\rm dS}=1$ and $c_{\rm AdS}=-2\sqrt{A_{\rm AdS}(A_{\rm AdS}-A_{\rm dS})}=-2$ , which corresponds to the values of the physical parameters $\kappa\tilde{V}_{0}\simeq 1.37$, $\kappa\tilde{V}_{\rm AdS}=-1.21$, and the derived parameter $\kappa\tilde{V}_{min}\simeq-8.97$.
@@
where the lower bound is the BF-bound Breitenlohner and Freedman (1982a, b). Moreover, since $\Delta_{-}=1<d/2=3/2$ , we are in a situation where we can choose multiple different quantizations for the scalar field, since both modes are normalizable. In particular if we expand it around infinity we find that
@@
Two typical choices correspond to setting the coefficient of the $\Delta_{-}$ term to be the source and that of $\Delta_{+}$ to be the VEV (Dirichlet BCs), and the other is the case in which $\tilde{\phi}_{-}$ corresponds to the VEV and $\tilde{\phi}_{+}$ to the source (Neumann BCs) [^17] In general one can consider mixed BCs leading to generic multi-trace deformations of the boundary CFT.. In addition, we observe that we are currently in a resonant setup, in the sense that we also have $\Delta_{+}-\Delta_{-}=1$ , so that the first subleading term in the expansion of $\tilde{\phi}_{-}$ scales the same as the leading term in the expansion of $\tilde{\phi}_{+}$ and so forth [^18] Nevertheless, no logarithmic term (associated with a conformal anomaly) appears in the expansion, since the conformal dimension satisfies $\Delta_{+}\notin\frac{d}{2}+\mathbb{N}$ (see the criterion in Petkou and Skenderis (1999))..
@@
We then write down the scalar counterterm action (for a single boundary component) [^19] This counterterm action is the same for all different choices of quantization. In general it is derived integrating the relation $\delta S_{c.t.}(\tilde{\phi})=\int_{\partial\mathcal{M}}d^{3}x\sqrt{h}\sum_{n<\Delta_{+}}\tilde{\pi}_{n}\delta\tilde{\phi}$ , where $\tilde{\pi}_{n}$ are the various terms in the expansion of the momentum., see Papadimitriou (2007) and references within
@@
In our specific case where $\Delta_{+}=2,\,\Delta_{-}=1<3/2$, only the first term in this action is needed for renormalization Papadimitriou (2007). We then compute the counterterm contribution to the momentum from (70) to be $\tilde{\Pi}_{c.t.}=\sqrt{h}\Delta_{-}H_{\rm AdS}\tilde{\phi}$. Subsequently, the renormalized momentum is given by the sum of the bare momentum (69) and the counterterm $\tilde{\Pi}_{c.t.}$ i.e.
@@ 4.3.2 The wineglass E/M gauge field background
Furthermore, if we wish to understand the wormhole background as providing good initial conditions for phenomenologically realistic inflation as in the scenario proposed in Betzios and Papadoulaki (2024a); Betzios et al. (2025), we can constrain the parameters of the model further. In particular assuming that at $\tau=t=0$, i.e. at the onset of standard Lorentzian evolution, the field is already located in the plateau region we obtain that,
$\tilde{V}_{0}\lesssim 3.73\times 10^{-10}\kappa^{-1}\left(\frac{r_{b}}{0.036}\right)\,$, (86)
where $r$ is the tensor-to-scalar ratio, bounded by $r<r_{b}=0.036$ at $95\%$ CL when combining the latest Planck, BICEP/Keck, and BAO data Akrami and others (2020); Ade and others (2021). Equations (56) and (57) allow one to relate the value of the scalar potential at $\tau=0$ to that of the E/M sector, namely
@@
$|H_{0,WG}|\gtrsim 3.17\times 10^{4}\left(\frac{0.036}{r_{b}}\right)^{1/2}\,.$ (88)
@@ 4.4.1 Piecewise oscillatory wormhole solutions
These results for the (quasi)-oscillatory wormholes are very important and contain crucial differences and resolve some paradoxes that appear in analogous (exactly) periodic solutions, such as the axion de-Sitter wormholes of Halliwell and Myers (1989); Aguilar-Gutierrez et al. (2023). The axion de-Sitter wormholes exist for a fixed cosmological constant, have no UV AdS boundaries and give rise to a serious paradox, since their number of periods is not physically bounded by any principle. Moreover, in that case each period lowers the on-shell action, rendering it unbounded from below in the limit of infinite periods [^22] In a sense this issue is related to the well-known conformal mode problem of Euclidean gravity, exacerbated by the shift symmetry of the constant potential (due to the fixed cosmological constant), see also Blommaert et al. (2025).. Here the presence of the AdS boundaries and the fact that the potential is a non trivial (quasi)-periodic function whose number of periods is directly correlated to the number of geometric periods in the IR throat part of the geometry, precludes such issues.